# Statistical Taylor Expansion: A New and Path-Independent Method for Uncertainty Analysis 111

### v17

@@ 2.6 Statistical Bounding
Figure 4: Measured bounding leakage $\epsilon(\kappa_{s},N)$ (y-axis) for varying measuring bounding range $\kappa_{s}$ (x-axis) and sample count NN (legend).
Figure 5: Measured bounding range $\kappa$ (left y-axis) and corresponding measured bounding leakage $\epsilon(\kappa)$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is uniform or normal (legend), with different measuring bounding range $\kappa_{s}$ for the normal distribution.
When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approach the distribution mean $\mu$ and distribution deviation $\sigma$ respectively as the sample count $N$ increases [4]. This yields the sample bounding leakage $\epsilon(\kappa,N)$ for the interval $\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$, in contrast to the distributional bounding leakage $\epsilon(\kappa)$ for the interval $\mu-\varrho\sigma,\mu+\kappa\sigma]$. Because $\epsilon(\kappa)\neq\epsilon(\kappa,N)$ for finite $N$, let $\epsilon(\kappa)=\epsilon(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bonding range, and $\kappa(\kappa_{s},N)$ is the measured bounding range.
@@
When the underlying distribution is Normal, Formula (2.28) and (2.29) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 4 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction. It also shows that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 5 further demonstrates that for smaller $\kappa_{s}$, $\kappa(\kappa_{s},N)$ approaches $\kappa_{s}$ more rapidly as $N$ increases, but converges to a larger stable bounding leakage $\epsilon(\kappa_{s})$.
When the underlying distribution is uniform, the portion of sampled range $\overline{x}-\sqrt{3}\delta x,\overline{x}+\sqrt{3}\delta x]$ outside the actual range $\mu-\sqrt{3}\sigma,\mu+\sqrt{3}\sigma]$ contributes to the bounding leakage $\epsilon(N)$. Figure 5 shows that $0<\epsilon(N)\sim N^{-0.564}$ empirically. The measured bounding range $\kappa(N)=\sqrt{3}(1-\epsilon(N))$.
2.7 Ideal Statistics
Figure 6: The ratio of resulting variance $\delta^{2}f$ to the ideal variance $\widehat{\delta^{2}}f$ (left y-axis) and the bounding leakage (right y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when the uncertainty distribution is Gaussian.
Figure 7: The ratio of resulting variance $\delta^{2}f$ to the ideal variance $\widehat{\delta^{2}}f$ (left y-axis) and the bounding leakage (right y-axis) for varying sample count NN (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when the uncertainty distribution is Uniform.
$\zeta(2,\kappa)=$ $\frac{\delta^{2}x}{(\delta x)^{2}};$ (2.30)
$\alpha(x\pm y)=$ $\frac{\zeta_{x}(2,\kappa_{x})(\delta x)^{2}+\zeta_{y}(2,\kappa_{y})(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.31)
When sample count $N\rightarrow\infty$, $\zeta(0)\rightarrow 1$, $\zeta(2)\rightarrow 1$ and $\delta^{2}x\rightarrow(\delta x)^{2}$ in Formula (2.30). Such relationship of $\delta^{2}f$ versus sample count $N$ is general:
  • When the underlying uncertainty distribution is Gaussian and $\kappa_{s}=5$, Figure 6 shows that the resulting variances $\delta^{2}f$ for a selected functions rise with $N$ until reaching the corresponding stable values $\widehat{\delta^{2}}f$ when $N\geq 50$.
  • When the underlying uncertainty distribution is Uniform, $\kappa_{s}=\sqrt{3}$, Figure 7 shows the same trend, however the stable variances $\widehat{\delta^{2}}f$ are reached only when $N\geq 10^{4}$.
The trends of uncertainty biases $\overline{f}-f$ versus sample count $N$ are exactly as those in Figure 6 and 7 respectively, except that $\overline{x}-x=0$. This similarity is expected because the first-order approximation in both Formula (2.5) and Formula (2.3) contains $\sigma^{2}\zeta(2,\kappa)$.
In both Figure 6 and 7, the bonding leakage $\epsilon(N)$ decreases as the sample count $N$ increases. The stable resulting variances $\widehat{\delta^{2}}f$ is reached only when $\epsilon(s,N)<10^{-3}$. All functions approach their corresponding stable variances with exactly the same trend in each figure. These observations suggest that bounding leakage is the reason for $\delta^{2}f<\widehat{\delta^{2}}f$ and it is justified to infer the confidential interval range $\kappa$ from bounding leakage $\epsilon(\kappa_{s})=1-\zeta(0,\kappa_{s})$ in Figure 4, or more precisely $1-\zeta(2,\kappa_{s})$ for $\delta^{2}f$.
Define the stable variance when the sample count $N$ is sufficiently large as the ideal variance $\widehat{\delta^{2}}f$. The choice of $\kappa_{s}$ to calculate $\widehat{\delta^{2}}f$ is defined as the ideal bounding range $\hat{\kappa}$. For Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$. When the input distribution is unbounded, the choice of $\hat{\kappa}$ is a trade-off: a larger $\hat{\kappa}$ means a smaller bounding leakage for the result but a narrower convergence range for the input. For Gaussian distribution, according to the 5-$\sigma$ rule for determining the statistical significance of experimental result [1][2][4], $\hat{\kappa}=5$.
@@ 2.7 Ideal Statistics
When the sample count $N$ is insufficient, $\widehat{\delta^{2}}f$ is calculated using $\hat{\kappa}$, to assume that $N$ is sufficient. $\delta^{2}f$ is calculated with the actual $N$. Define the ideal ratio as $\alpha\equiv\delta^{2}f/\widehat{\delta^{2}}f$, which is 0 when $N=1$ and $1$ when $N$ is sufficiently large. The ideal ratio $\alpha$ quantifies the reliability of $\widehat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.31) as an example for $\alpha(x\pm y)$ .
@@ 2.8 Dependency Tracing
  • The dependency tracing of $\delta^{2}f(g(x))$ is demonstrated by $\delta^{2}(f^{-1}(f(x)))=(\delta x)^{2}$. For a reversible transformation such as matrix reversion or FFT, after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original input should be restored.
@@ 2.9 Traditional Execution and Dependency Problem
Dependency tracing requires an analytic form of the function to apply statistical Taylor expansion for the result mean and variance, as in Formula (2.5), (2.3), (2.8), and (2.3). This requirement often conflicts with conventional numerical methods for analytic functions:
@@
  • Furthermore, traditional approaches frequently apply approximations to result values during execution. Under the statistical Taylor expansion, Formula (2.3) shows that the variance converges more slowly than value in statistical Taylor expansion. Consequently, approximation strategies should prioritize variances than values.
  • Traditionally, results from mathematical library functions are accepted without scrutiny, with accuracy assumed down to the last digit. As demonstrated in Section 10, statistical Taylor expansion enables the detection of numerical errors within these functions and requires that they be recalculated with uncertainty deviations explicitly incorporated into the output.
  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly and independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls. However, this decomposition introduces dependency problems. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
@@ 3 Variance Arithmetic
Because of the finite precision and limited range of conventional floating-point representation, $\zeta(n)$ can only be computed to limited terms. Consequently, the following numerical rules are introduced:
@@
For simplicity of discussion, the Taylor coefficients in Formula (2.4) and (2.7) are assumed to be precise.
@@ 3.1 Monotonic
Figure 8: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for Gaussian uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$. When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
Figure 9: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for uniform uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). The x-axis is expressed in units of $\pi$. When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure. When $c\in[0,1]$, uncertainty bias is negative so it does not show up in the plot.
Beyond an upper bound $\delta x$, the expansion is no longer monotonic for $e^{x\pm\delta x}$, $\log(x\pm\delta x)$, and $(x\pm\delta x)^{c}$.
For Gaussian input uncertainty with $\hat{\kappa}=5$:
@@
  • For $(x\pm\delta x)^{c}$, except when $c$ is a natural number, the upper bound $P(x)$ is close to $1/5$ but increasing with $c$. This trend is displayed in Figure 8.
Similar trends holds when the input uncertainty is uniform but with different values. For example, after the respective upper bound $P(x)$ for $(x\pm\delta x)^{c}$ are normalized by the corresponding $\hat{\kappa}$, they show almost the same trends of increasing with $c$. Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$.
@@ 3.2 Positive
Besides convergence, the variance expansion may yield negative results, as in Formula (2.20) for $sin(x\pm\delta x)$. Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty. Beyond this upper bound, the result deviation is no longer positive at some expansion order. Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which Figure 10 satisfies. Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
@@ 3.3 Floating-Point Rounding Errors
Variance arithmetic incorporates floating-point rounding errors as the $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$. Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP of $x$, where ULP refers to the Unit in the Last Place in conventional floating-point representation [8], because rounding errors are shown to be uniformly distributed within ULP [17].
@@ 4 Verification of Variance Arithmetic
Once input error from every source is accounted precisely, ideal coverage is achieved in either contexts:
  • Distribution Test: When comparing the calculated mean and deviation with the result data set, the error deviation is exactly $1$ and the error distribution is Normal. Empirically, error distributions are close to Normal when error deviation are close to 1, regardless of input uncertainty distribution, and such convergence occurs rapidly [17]. This is due to central limit theorem [4].
  • Value Test: When comparing the calculated value and the result value in a corresponding data set, the error deviation is much less than $1$ and the error distribution is Delta. For example, a round-trip test is a value test.
@@ 5 Polynomial
Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can only reach half of the maximal expansion order of Formula (2.3), for example, $224=448/2$ when the input uncertainty is assumed to be Gaussian.
@@ 5.1 Residual Error
Figure 12: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x-axis). The y-axis to the left shows both the value and the uncertainty of the residual errors. The y-axis to the right indicates the expansion order needed to reach stable value for each $x$.
Figure 12 shows the residual error of $\sum_{j=0}^{224}x^{j}-1(1-x)$. It also displays the required expansion orders for 1/(1x)1/(1-x), which are all less than $224$. Therefore, the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$. A detailed analysis indicates that the maximal residual error is $4$ times the ULP of 1/(1x)1/(1-x). In all cases, the calculated uncertainty bounds the residual error effectively for all $x$, with an error deviation of $2.60$ when the expansion order is less than $224$. Variance arithmetic can provide proper coverage for rounding errors.
@@ 5.2 Continuity
In variance arithmetic, the result mean, variance and histogram are generally continuous across parameter space. For example, $\delta x$ has an upper bound for $(x\pm\delta x)^{c}$ to converge except when $c$ is a natural number $n$. The result mean, variance and histogram of $(x\pm\delta x)^{c}$ remain continuous around $c=n$.
5.3 Distributional Hole
Figure 13: Histograms of normalized errors for $(x\pm 0.2)^{n}$, with x=0,0.2,+0.2x=0,-0.2,+0.2, and $n=2,3$, as indicated in the legend.
Unlike a distributional zero, a statistical bounding range in variance arithmetic can include a distributional pole, such as around $(0\pm\delta x)^{c},c>1$. The presence of such poles does not disrupt the continuity of the result mean, variance, or histogram. Figure 13 illustrates the error distributions of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.

### v18

@@ 2.6 Statistical Bounding
Figure 4: Measured linear leakage $\omega(\kappa_{s},N)$ (y-axis) for varying measuring bound range $\kappa_{s}$ (x-axis) and sample count NN (legend).
Figure 5: Measured linear leakage $\omega(\kappa,N)$ (left y-axis) and corresponding measured bound range $\kappa$ (right y-axis) for varying sample count NN (x-axis) when the underlying distribution is Gaussian, with different measuring bound range $\kappa_{s}$ (legend).
Figure 6: Ideal leakage (y-axis) for varying sample count $\kappa_{s}$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=5$ for Gaussian input uncertainties.
Figure 7: Ideal leakage (y-axis) for varying bound range $\kappa$ (x-axis) for the selected function $f(x=1\pm 0.1)$ (legend) when $\hat{\kappa}=6$ for Gaussian input uncertainties.
$f(x)-\overline{f(x)}$ $\simeq(\delta x)^{2}\zeta(2,\kappa)f_{x}^{(2)};$ (2.28)
$\delta^{2}f(x)$ $\simeq(\delta x)^{2}\zeta(2,\kappa)(f_{x}^{(1)})^{2};$ (2.29)
Figures (2.28) and (2.29) show that the first-order approximation of Formulas (2.5) and (2.3) both contain the term $(\delta x)^{2}\zeta(2,\kappa)$. As $\kappa\rightarrow+\infty$, $\zeta(2,\kappa)\rightarrow 1$ such that $\delta^{2}f(x)$ reaches a stable value but the convergence range of $\delta^{2}f(x)$ reduces toward zero for Formulas (2.17) and (2.22). As a trade-off, the choice of $\kappa$ to calculate the stable $f(x)-\overline{f(x)}$ and $\delta^{2}f(x)$ is defined as the ideal bound range $\hat{\kappa}$, and the result is the ideal uncertainty bias $\widehat{\overline{f(x)}}-f(x)$ and ideal variance $\hat{\delta^{2}}f(x)$. For a Uniform distribution, by definition $\hat{\kappa}=\sqrt{3}$. For a Gaussian distribution, according to the 5-$\sigma$ rule for determining the statistical significance of experimental result [1][2][4], $\hat{\kappa}=5$ by default.
Define linear leakage $\omega(\kappa)\equiv 1-\zeta(2,\kappa)$. When sampling from a distribution, the sample mean $\overline{x}$ and sample deviation $\delta x$ approach the distribution mean $\mu$ and distribution deviation $\sigma$ respectively as the sample count $N$ increases [4]. This yields the sample linear leakage $\omega(\kappa,N)$ for the interval $\overline{x}-\varrho\delta x,\overline{x}+\kappa\delta x]$, in contrast to the distributional linear leakage $\omega(\kappa)$ for the interval $\mu-\varrho\sigma,\mu+\kappa\sigma]$. Because $\omega(\kappa)\neq\omega(\kappa,N)$ for finite $N$, let $\omega(\kappa)=\omega(\kappa_{s},N)$, where $\kappa_{s}$ is the measuring bound range, and $\kappa(\kappa_{s},N)$ is the measured bound range, with the latter used in calculating $\zeta(n,\kappa)$.
$\epsilon(\kappa)$ $=\omega(\kappa)+2N(\kappa)\kappa;$ (2.30)
@@
Define bounding leakage $\epsilon(\kappa)\equiv 1-\zeta(0,\kappa)$. When the underlying distribution is Normal, Formula (2.30) presents the difference between the linear leakage $\omega(\kappa)$ and the bounding leakage $\epsilon(\kappa)$, while Formulas (2.31) and (2.32) give the distributional bounding leakage $\epsilon(\kappa)$ and the sample bounding leakage $\epsilon(\kappa_{s},N)$ respectively, where $\xi()$ is the Normal error function [4]. Figure 4 shows that $\epsilon(\kappa_{s})<\epsilon(\kappa_{s},N)$, and $\lim_{N\rightarrow\infty}\epsilon(\kappa_{s},N)=\epsilon(\kappa_{s})$ along the y-axis direction, and that $\kappa(\kappa_{s},N)<\kappa_{s}$ and $\lim_{N\rightarrow\infty}\kappa(\kappa_{s},N)=\kappa_{s}$ along the x-axis direction. Figure 5 sections Figure 4 along the y-axis direction for $\kappa_{s}=4,5,6$, and shows that to reach the stable variance for a $\kappa_{s}$, the required sample count is $N\gtrsim 10^{3}$, whereas the stable result bounding leakage is smaller for larger bounding range $\kappa$. It also shows that linear leakage converges to 0 with increasing sample count $N$ faster when $\hat{\kappa}$ is larger.
$\varepsilon(x\pm y)$ $=\frac{\omega_{x}(\delta x)^{2}+\omega_{y}(\delta y)^{2}}{(\delta x)^{2}+(\delta y)^{2}};$ (2.33)
@@
Define ideal leakage as $\varepsilon\equiv 1-\delta^{2}f/\widehat{\delta^{2}}f\simeq\omega$, which quantifies the reliability of $\hat{\delta^{2}}f$ from the underlying distributions and the corresponding sample counts of all inputs, with Formula (2.33) as an example. Figure 6 shows that the stable ideal leakages for the selected functions are less than $10^{-4}$ for $\hat{\delta^{2}}f$, which is small enough for most applications. Figure 7 compares the ideal leakages for $\hat{\delta^{2}}f$ for the selected functions when different $\hat{\kappa}$ is used.
@@ 2.7 Dependency Tracing
  • The dependency tracing of $\delta^{2}f(g(x))$ is demonstrated by $\delta^{2}(f^{-1}(f(x)))=(\delta x)^{2}$. For a reversible transformation such as matrix reversion or FFT (Fast Fourier Transformation), after a round-trip transformation which is a forward transformation followed by a reverse transformation, the original input should be restored.
@@ 2.8 Traditional Execution and Dependency Problem
Dependency tracing requires applying statistical Taylor expansion on an analytic form of the function in one step. This requirement often conflicts with conventional numerical methods for analytic functions:
@@
  • In conventional practice, an analytic expression is often decomposed into simpler, ostensibly independent arithmetic operations such as negation, addition, multiplication, division, square root, and library calls. However, this decomposition introduces dependency problems. For example, if $x^{2}-x$ is calculated as $x^{2}-x$, x(x1)x(x-1), and $(x-\frac{1}{2})^{2}-\frac{1}{4}$, only $(x-\frac{1}{2})^{2}-\frac{1}{4}$ gives the correct result, while the other two give wrong results for wrong independence assumptions between $x^{2}$ and $x$, or between x1x-1 and $x$, respectively.
@@ 3 Variance Arithmetic
Because of the finite precision and limited range of the floating-point representation, $\zeta(n,\kappa\leq 5)$ can be computed only to limited terms. Consequently, the following numerical rules are introduced:
@@
For simplicity of discussion, the Taylor coefficients in Formulas (2.4) and (2.7) are assumed to be precise.
@@ 3.1 Monotonic
Figure 8: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for Gaussian uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure.
Figure 9: Measured upper bound $\delta x$ (left y-axis) for $(1\pm\delta x)^{c}$ across different values of $c$ (x-axis) for uniform uncertainty. The corresponding uncertainty bias and uncertainty are also shown (right y-axis). When $c$ is a natural number, $\delta x$ has no upper bound; however, such cases are omitted in the figure. When $c\in[0,1]$, uncertainty bias is negative so it does not show up in the plot.
Beyond an upper bound $\delta x$, the expansion is no longer monotonic for $e^{x\pm\delta x}$, $\log(x\pm\delta x)$, and $(x\pm\delta x)^{c}$. For Gaussian input uncertainty with $\hat{\kappa}=5$:
@@
  • For $(x\pm\delta x)^{c}$, except when $c$ is a natural number, the upper bound $P(x)$ is close to $1/5$ but increasing with $c$. This trend is shown in Figure 8.
Similar trends holds when the input uncertainty follows other distribution. Figure 9 shows the upper bound $P(x)$ for $(x\pm\delta x)^{c}$ when the input uncertainty distribution is Uniform. It has almost the same trend of upper bounds $P(x)$ increasing with $c$ as shown in Figure 9 but just with different y-scaling: $1/\sqrt{3}$ instead of by $1/5$.
@@ 3.2 Positive
In addition to convergence, the variance expansion may yield negative results, as in Formula (2.20) for $sin(x\pm\delta x)$. Figure 10 shows that the upper bound of $\delta x$ for $sin(x\pm\delta x)$ varies periodically between $0.318\pi$ and $0.416\pi$ for Gaussian input uncertainty. Beyond this upper bound, the result deviation is no longer positive at some expansion order. Conceptually, the upper-bound for $\delta x$ should not be much more than $\pi/2$ to make the result meaningful, which is satisfied as shown in Figure 10. Similar trend holds when the input uncertainty is uniform but with larger upper bound $\delta x$, as shown in Figure 11.
@@ 3.3 Floating-Point Rounding Errors
Variance arithmetic incorporates floating-point rounding errors as the $\delta x$ when converting a floating-point value $x$ into $x\pm\delta x$. Unless all the least 20 bits of the significand of $x$ are zero, $\delta x$ is assumed to be $1/\sqrt{3}$ times of the ULP (Unit in the Last Place [8]) of $x$, because rounding errors are shown to be uniformly distributed within ULP [17].
@@ 4 Verification of Variance Arithmetic
Once input error from every source is accounted precisely, ideal coverage is achieved in either context:
  • Distribution Test: When comparing the calculated mean and deviation with the result data set, the error deviation is exactly $1$ and the error distribution is Normal, regardless of input uncertainty distribution. Such convergence to Normal distribution occurs rapidly [17] because of the central limit theorem [4].
  • Value Test: When comparing values one-by-one between a calculated data set and the corresponding result data set, the error deviation is much less than $1$ and the error distribution is Delta. For example, a round-trip test is a value test.
@@ 5 Polynomial
Because the maximal expansion term using Formula (5.1) is $\tilde{x}^{2N}$, $N$ in Formula (5.1) can reach only half of the maximal expansion order of Formula (2.3), for example, $224=448/2$ when the input uncertainty is assumed to be Gaussian, and $\hat{\kappa}=5$.
@@ 5.1 Residual Error
Figure 12: Residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$ vs $x$ (x-axis). The left y-axis shows both the value and the uncertainty of the residual errors. The right y-axis indicates the expansion order needed to reach stable value for each $x$.
Figure 12 shows the residual error of $\sum_{j=0}^{224}x^{j}-\frac{1}{1-x}$. It also displays the required expansion orders for $\frac{1}{1-x}$, which are all less than $224$. Therefore, the residual error reflects solely the rounding error between $\sum_{j=0}^{224}x^{j}$ and $\frac{1}{1-x}$. A detailed analysis indicates that the maximal residual error is four times the ULP of $\frac{1}{1-x}$. In all cases, the calculated uncertainty bounds the residual error effectively for all $x$, with an error deviation of $2.60$ when the expansion order is less than $224$. Variance arithmetic can provide proper coverage for rounding errors.
@@ 5.2 Continuity
In variance arithmetic, the result mean, variance and error distribution are generally continuous across parameter space. For example, when $c$ is a natural number $n$, $(x\pm\delta x)^{c}$ becomes a polynomial which has no upper bound for $\delta x$, in contrast to when $c$ is not a natural number as shown in Figure (8). However, the result mean, variance and error distribution of $(x\pm\delta x)^{c}$ remain continuous across $c=n$.
5.3 Distributional Pole
Figure 13: Error distributions for $(x\pm 0.2)^{n}$, with x=0,0.2,+0.2x=0,-0.2,+0.2, and $n=2,3$, as indicated in the legend.
A statistical bounding range in variance arithmetic can include a distributional pole, such as around $(0\pm\delta x)^{c},c>1$. The presence of such poles does not disrupt the continuity of the result mean, variance, or error distribution. Figure 13 illustrates the error distributions of $(x\pm 0.2)^{n}$ when x=0,0.2,+0.2x=0,-0.2,+0.2 and $n=2,3$.